FEL Upgrade Project Weekly Brief
February 12-16, 2001







Highlights:

We operated the IR Demo successfully for four (4) users and continued to test prototypes of systems that will be used in the Upgrade.

We successfully ran all eight (8) klystrons in Zone 4 at 5 kW simultaneously.

Management:

We worked on revising Phase I cost performance plan to include the engineering change request approved at the October review and also worked on developing a cost performance plan for the Phase II activities which will start on or about March 1.

The Upgrade Program Review Committee has chosen March 27 and 28 for the next semi-annual review of the Upgrade Program. The first day of the review will be a status report by each WBS Manager.  The second day of review will be a workshop on scaling issues that will include contributions from the members of the review panel.

We have several high level visits scheduled for the month of March for briefings and tours on the FEL Project:
    1)    March 1    John Eric from the Air Force Research Lab
    2)    March 2    DOD Joint Technology Office
    3)    March 22    Adms Mathis and Nanos
 

On Feb. 15 we participated in a workshop sponsored by the NSF and the ODU School of Engineering.  ODU is a finalist in the competition for an Industry University Cooperative Research Center (IUCRC) to be hosted in the Applied Research Center. The proposed center would focus on laser processing.

WBS 3 (Beam Physics):

A detailed UV bypass Revision 1.1 layout for transport from linac to wiggler was developed. Initial study suggests it will provide adequate performance at a wiggler located in the pit. Subsequent work will address the layout and performance of the bypass energy recovery.

WBS 5 (SRF):

Prototype tuner #2 test and evaluation is underway.
Received the deep draw fixture for the fundamental power coupler body and QA looks good.
Question back out to thermal shield vendors from initial proposals.
Detailing cold valve details with vendor for quote.
 

WBS 6 (RF):

Zone 3 - No progress.

Zone 4 - All 8 klystrons have been run at 5 kW simultaneously.  The calibration tests have not yet been run due to incomplete calibration software.  This software is in progress.  A "punch list" of 12 items needs to be completed before the zone is finished.  The most serious is a Cathode Power Supply that trips twice each morning and then is fine for the rest of the day.

Quarter HVPS - No progress with the vendor PSC.  Their engineer was traveling this week.  Discussions are planned for next week. The 100 kW klystrons by CPI are on schedule.  They plan to ship the first one on April 18, 2001.

Gun HVPS - Plans are being made to extend the length of the Gun's HV Tank.  An additional multiplier stack must be added for the additional Gun current.  It appears the Gun Tank drawings were never released, but I have found many of the drawings that were used to build it.

WBS 8 (I&C):

The new surface mount Optical Beam Position Monitors (OBPMís) have been received and are being evaluated; there are mechanical problems that require new boards to be manufactured. The sample and hold delay circuit was optimized for the OBPM system to allow for better measurement of beam position and to allow better positioning of monitors. The new OPBM/optical transport picomotor control chassis is complete and ready for testing. The cable for the commercial quad-detectors was tested for Optics.

Drawings have come back from EECAD for GPIB System drawing and an update to the LSS User Interface Box in the user labs and should be signed off soon.  New drawings for the remote shutter control boxes and the remote mirror control box have been submitted.

The first article (2 pcs) of the 3" shielded beam viewer are in an being tested; they look great and have passed leak detection. The residual gas analyzer (RGA) output showed  no contaminates. The 3" BPMs are due in a week.

WBS 9 (Transport):

Dipoles
Injector Dipoles (DU/DV)
o  No report from DULY Research
Arc Dipoles (GY, GX, GQ)
o  No report from AES
Optical Chicane Dipole (DW)
o  Revisions to the drawing package are being made by two designers on a part-time effort.  Roughly 25%
    of the revisions have been completed.
Quadrupoles
QX (3.125" Quad)
o Measurement Probe: final machining of the coil support plate is progressing with an expected delivery
    early next week.
o Revision of the drawing package is progressing.
Sextupole with H/V Correctors
o Tosca/2D model is being fine-tuned.
BT Layout
o Layouts are in progress to optimize the use of single and/or double quad girders.

WBS 10 (Wiggler):

The dispersion section coils were received by PECO and assembly to the core has begun. A Radia model
for the optical klystron wiggler has been developed and is being debugged and checked out.  The initial field profile in the center of the wiggler agrees with previous calculations using Poisson.

Optical Klystron
o Coils were received by PECO and assembly to the core has begun.

WBS 11 (Optics):

We finished our review of the check prints for the deformable mirror assembly, and the drawing package is being released for bid. The design team continues to work on the upgrade optical transport system (OTS) and optical cavity assemblies.

We also supported operations.  Highlights were:

We had the first opportunity to see how well the hutch covering the optical diagnostics worked.  It contained the nitrogen atmosphere quite well, greatly reducing measurement problems in the 6 micron spectral region.  We built a test stand for testing the O-BPM detector assemblies that doesn't involve laser sources.  This makes testing more flexible.  With it, the O-BPMs in several locations were removed, checked, balanced, and, where appropriate, reinstalled.  We also installed neutral density filters
with less attenuation in the filter wheels before the detector assemblies.

There were some issues on drive laser output stability early in the week that we addressed.

Operations/Commissioning:

We had a good week of running the machine.  We spent several sessions looking at IR Microscopy using an atomic force microscope.  We obtained our first IR images of a surface using this technique.  It is expected to be able to provide sub-micron resolution on surface absorption in the mid-IR.  Initial experiments were at 5.85 microns.

We also did our first experiments on non-linear dielectric absorption and saw good signals at 18.7 MHz and excellent signal to noise ratios.  Those results are presently under analysis.

We provided several runs for carbon nanotube generation for a NASA/College of William and Mary collaboration at around 6 microns and some interesting results operating the laser with 30 millisecond macropulses.  Overall it is not clear that six micron region is optimal for this work, though.  An answer to that awaits analysis of the nanotube product.

On Friday we recesiated the cathode to provide more charge for next week's effort.  We also performed work on the optical beam position monitor system which had not been performing to expectations.  It was discovered that many of the detectors had significant dead areas which were interfering with accurate quadrature measurements.  We are examining alternatives.
 

Privacy and Security Notice
Privacy and Security Notice